OpenStudy (anonymous):

Solve for n:

OpenStudy (anonymous):

\[\frac{ 3n+2 }{ 7 } = 1- \frac{ 2n-3 }{ 4 }\]

OpenStudy (anonymous):

-1/26 41/26 10/13 7/13

OpenStudy (yttrium):

\[4(3n+2) = 28 - 7 (2n-3)\] \[12n+8 = 28 -14n +21\] \[26n = 41\] \[n = 41/26\]